2. Center-of-gravity (cg) location is critical for airplane stability. To get the empty og location, the Piper J-4 Cub pict2. Center-of-gravity (cg) location is critical for airplane stability. To get the empty cg location, the Piper J-4 Cub pict
Answer #1

Answer for point f is added at the end, please don't consider the point written in the middle.Answer! а) 16tet emps of ght = 436 + 444 45% total = 940 Lb b) Idealising the Shown below. 1 fuel Dotom tanley aircraft as an

taking moment about B = Ro (1829-3:26) aury 168.5et 350(1825-23) a RA 2 1196.42 loods can be The moments mts produces 2 SeperI . MAC in de 16:38 L267 1 h) al. In configuration : ď ca is Alt of datum 10 gallons a 83.5 l but following the same approach